OpenStudy (anonymous):

The measure of ∠A is 35° more than three times the measure of ∠B. The sum of their measures is 135°. Find m ∠A and m ∠B.

OpenStudy (wolf1728):

A = 35° B= 3*A = 105° A + B = 140° I get the sum as 140 Also, you want 'm' of each angle - do you mean slope?

OpenStudy (anonymous):

They want the sum of the measurements to equal 135 degrees

OpenStudy (wolf1728):

Well, if A is 35 and B is 3 times A then their sum will always be 140

OpenStudy (anonymous):

A isn't 35, A is 35 degrees MORE than 3 times the measure of B

OpenStudy (wolf1728):

Sorry I misread that. A = (3 * B) + 35 A + B =135 135 -B = 3B + 35 100 = 4B B = 25 A = 110

OpenStudy (anonymous):

Thanks!

OpenStudy (wolf1728):

thanks for the medal